# Calibrating and Rotating: A Unified Framework for Weight Conditioning in PEFT

### v1

@@ Introduction
Large-scale Pre-trained Models constitute the bedrock of modern artificial intelligence, yet their substantial fine-tuning costs drive research into Parameter-Efficient Fine-Tuning (PEFT) (Ding et al. 2023). PEFT techniques adapt these models with minimal computational and storage overhead through diverse strategies, ranging from inserting small, trainable adapter modules into the model’s architecture (Houlsby et al. 2019a) to methods like prompt-tuning, which only modify the input embeddings (Lester, Al-Rfou, and Constant 2021a; Li and Liang 2021). Among these, LoRA (Hu et al. 2021) is one of the most widely adopted PEFT techniques, which operates by adding a low-rank update $\mathbf{BA}$ to the pre-trained weights $\mathbf{W_{\text{pre}}}$ (i.e., $\mathbf{W=W_{pre}}+s\cdot\mathbf{BA}$).
The simplicity and effectiveness of LoRA inspire a continuous stream of research, leading to numerous variants that improve upon the original method along several key axes. These advancements primarily focus on: (i) dynamic rank allocation during training, as exemplified by AdaLoRA (Zhang et al. 2023) and ReLORA (Lialin et al. 2023); (ii) improved initialization of low-rank matrices from pre-trained weights, proposed in methods like PiSSA (Meng, Wang, and Zhang 2024) and LoRA-GA (Wang, Yu, and Li 2024); and (iii) refined optimization strategies, such as those introduced in LoRA+ (Hayou, Ghosh, and Yu 2024) and LoRA-RITE (Yen et al. 2024).
Our work connects to the thread of principled weight preparation, with a focus on DoRA (Weight-Decomposed Low-Rank Adaptation) (Liu et al. 2024). As an improvement upon LoRA, DoRA pushes this philosophy further by decoupling a pre-trained weight matrix $\mathbf{W}$ into its magnitude and direction. It then applies the low-rank update exclusively to the directional component, leaving the magnitude fixed. This process is formulated as $\mathbf{W}=\mathbf{m}\frac{\mathbf{W_{\text{pre}}}+s\mathbf{BA}}{\|\mathbf{W_{\text{pre}}}+s\mathbf{BA}\|_{c}}$.
@@
To answer this question, we conduct an in-depth empirical analysis. We find that while prior work points to stable rank (Lion et al. 2025), its explanatory power is limited in many practical tasks. Instead, we find that DoRA consistently increases the Singular Value Entropy of the weight update matrix. A higher entropy signifies a broader distribution of update energy across features, which prevents over-specialization and emulates the desirable behavior of full fine-tuning (Liu et al. 2025). This suggests DoRA’s success stems from indirectly optimizing the weight update’s singular value spectrum. To validate this and address DoRA’s inefficiency, we reframe it as a more efficient Weight Conditioning method (Saratchandran, Wang, and Lucey 2024). This perspective reveals DoRA’s mechanism as modulating the main weight matrix with a lightweight, learnable matrix. This discovery opens a broader design space, for which we propose a unified weight conditioning framework to create more powerful and efficient PEFT methods.
Within this framework, we introduce two novel methods: (i) Pre-Diag which efficiently calibrates pre-trained weights to improve performance while reducing training time, and (ii) Skewed Orthogonal Rotation Adaptation (SORA), a powerful, norm-preserving transformation that uses parameter-efficient orthogonal rotations to align the feature space, building on Orthogonal Fine-Tuning (OFT) (Qiu et al. 2023).
@@
To understand the performance discrepancies among different fine-tuning methods, we first analyze the stable rank of the parameter update matrix, $\Delta\mathbf{W}=\mathbf{W}-\mathbf{W}_{\text{pre}}$, where $\mathbf{W}$ represents the final weights and $\mathbf{W_{\text{pre}}}$ the initial pre-trained weights. The stable rank is defined as $\|\Delta\mathbf{W}\|_{F}^{2}/\|\Delta\mathbf{W}\|_{2}^{2}$ (Rudelson and Vershynin 2007). Our choice of this metric is motivated by prior work that has linked it to the diversity of update directions (Lion et al. 2025) and the expressive capacity of low-rank models (Zeng and Lee 2023; Kalajdzievski 2023). Our initial hypothesis was that the stable rank could quantitatively explain why methods like Full Fine-Tuning (FFT) and DoRA outperform LoRA.
@@
The limitations of stable rank motivate our search for a more discerning metric. We therefore employ singular value entropy, defined as the Shannon entropy of the normalized singular value spectrum: $H(\sigma)=-\sum_{i}p_{i}\log p_{i}$, where $p_{i}=\sigma_{i}^{2}/\sum_{j}\sigma_{j}^{2}$  (Roy and Vetterli 2007; Alter, Brown, and Botstein 2000). Intuitively, this metric measures how uniformly the update’s energy is distributed across its singular directions. A higher entropy indicates that the update is more uniformly distributed, rather than being dominated by a few high-magnitude singular values. This uniformity prevents the update from concentrating on only a few dominant features.
@@
This finding raises a crucial question: how does the DoRA formulation mechanically increase singular value entropy? The original DoRA update rule (Liu et al. 2024) decomposes the weight matrix $\mathbf{W}$ into a learnable magnitude $\mathbf{m}$ and a direction component:
@@
We can algebraically reformulate this operation as a form of weight conditioning (Saratchandran, Wang, and Lucey 2024). This reformulation reveals the effective weight change $\Delta\mathbf{W}_{\text{DoRA}}=\mathbf{W}-\mathbf{W_{\text{pre}}}$ to be:
@@
This higher-rank component is key. For singular value indices $i>r$, the LoRA update is zero (i.e., $\sigma_{i}(\Delta\mathbf{W}_{\text{LoRA}})=0$). In contrast, Weyl’s theorem (Horn and Johnson 2012) shows that DoRA maintains non-zero singular values in these dimensions:
@@
LoRA (Hu et al. 2021) $\mathbf{W_{\text{pre}}+BA}$ Identity Matrix N/A Low-Rank Adaption
DoRA (Liu et al. 2024) $\mathbf{(W_{\text{pre}}+BA)D}$ Diagonal Matrix Post Weight-Decomposed
@@
This approximation makes $\mathbf{P}$ near-orthogonal while drastically reducing computational demands. (Singla and Feizi 2021) first introduced the skew-orthogonal strategy into CNNs and provided a bounding for the approximation of matrix exponentials. Similarly, in SORA, for the orthogonal matrix $\mathbf{P}$, we have the following theorem:
@@
To test general language understanding, we fine-tuned DeBERTaV3-Base (He et al. 2020) on the GLUE benchmark(Wang et al. 2018). We test three learning rates ($\{1\text{e-}3,8\text{e-}4,4\text{e-}4\}$), each with three random seeds for robustness. Table 2 shows aggregated results comparing our Pre-Diag and SORA with LoRA and DoRA, reporting the best average performance. The findings indicate that our methods achieve superior average performance on the benchmark.
@@
For commonsense reasoning, we conduct several experiments. First, we train LLaMA3-8B (Dubey et al. 2024) on the Commonsense15k dataset and test its performance on a commonsense benchmark. The experimental results are presented in Table 3. SORA achieves the optimal average score across all tasks with different ranks, which effectively confirms the effectiveness of SORA.
@@
For mathematical reasoning, we conduct several experiments. First, we train Gemma-7B (Gemma Team et al. 2024) on the MetaMathQA14k dataset (Yu et al. 2023) and test its performance on a mathematical benchmark. The experimental results are presented in Table 4. SORA achieves the optimal average score across all tasks, which effectively confirms the effectiveness of SORA.
@@
Parameter-Efficient Fine-Tuning (PEFT) methods reduce the significant computational and storage costs associated with adapting large pre-trained models. Among various approaches, such as adapter modules (Houlsby et al. 2019b) and prompt-tuning (Lester, Al-Rfou, and Constant 2021b), methods based on low-rank adaptation are particularly prominent for their effectiveness and efficiency.
@@
LoRA (Hu et al. 2021) is a cornerstone of this approach. It approximates weight updates as the product of two low-rank matrices, based on the assumption that these updates occupy a low-dimensional manifold. This method is also provably capable of approximating any target transformer model under mild assumptions (Zeng and Lee 2023). This elegant mechanism inspires a wave of research that unfolds along three main threads.
@@
The first thread focuses on expressiveness and optimisation. Methods like AdaLoRA (Zhang et al. 2023) and ReLoRA (Lialin et al. 2023) replace LoRA’s static rank with dynamic allocation schedules to place capacity where it is most needed. Others enrich the update structure beyond a simple matrix product; for instance, BTT (Potapczynski et al. 2024) and CoMERA (Yang et al. 2024) introduce tensor factorisations while maintaining linear complexity. The training process itself is also refined. LoRA+ (Hayou, Ghosh, and Yu 2024) advocates for distinct learning rates for the two factors, while LORO (Zhanfeng Mo 2025) and ScaleAdam (Zhang and Pilanci 2024) reframe the optimisation on Riemannian manifolds to align updates with the parameter space’s intrinsic curvature.
@@
A second thread pursues efficiency under tight resource constraints. QLoRA (Dettmers et al. 2023) enables fine-tuning on a single consumer GPU by quantising the model backbone to 4-bits while keeping LoRA activations in higher precision. Complementary efforts further shrink memory footprints by compressing the adapters or sharing low-rank factors across layers and tasks, as seen in methods like LoQT and VeRA (Loeschcke et al. 2024; Kopiczko, Blankevoort, and Asano 2023).
@@
The third thread involves the principled initialisation of weights. PiSSA (Meng, Wang, and Zhang 2024) and OLoRA (Büyükakyüz 2024) initialise the low-rank factors using singular vectors or orthogonal bases from the pre-trained weights, aligning the initial update with salient directions. DoRA (Liu et al. 2024) extends this by decomposing each weight into magnitude and direction, applying LoRA only to the directional component. Such targeted adaptation yields significant gains and highlights the value of structuring the update, not just constraining its rank.
@@
Beyond LoRA-centric innovations, other structured PEFT methods explore different constraints. One class of methods, including HRA (Yuan, Liu, and Xu 2024) and HOFT (Arcas et al. 2025), constructs complex orthogonal matrices from simple low-rank operations like Householder transformations. Another evolutionary line of structured methods seeks to balance expressive power, cost, and efficiency. This line evolves from the restrictive block-diagonal form of OFT (Qiu et al. 2023), to the denser but more expensive butterfly structure of BOFT, and finally to the efficient group-and-shuffle structure of GSOFT (Gorbunov et al. 2024). Our method, SORA, also operates in this space. Inspired by the Cayley transform (Trockman and Kolter 2021), SORA constructs orthogonal matrices using a first-order Taylor expansion of the matrix exponential, an approximation also explored in (Singla and Feizi 2021). As a future direction, SORA can be integrated with orthogonalisation techniques to enhance performance.
@@ Acknowledgments
This research was supported by the Guangdong S&T Program (No. 2024B0101010003). The computing resources were provided by Pengcheng Cloud Brain.
@@ Appendix B Appendix 2: Proof of Theorem 2
See the proof of [(Singla and Feizi 2021), Theorem 3].
@@
where $(\circ)$ is due to the use of the triangle inequality; $(\star)$ is due to $\|\mathbf{A}\mathbf{B}\|_{F}\leq\|\mathbf{A}\|_{F}\|\mathbf{B}\|_{F}$ for all $\mathbf{A},\mathbf{B}\in\mathbb{R}^{m\times n}$.
@@ Appendix C Appendix 3: Experimental Details
We fine-tune the pre-trained DeBERTa-Base model(He et al. 2020) on the GLUE benchmark using the Hugging Face implementation [^1] https://huggingface.co/transformers/model˙doc/roberta.html [^2] https://huggingface.co/datasets/nyu-mll/glue. For all tasks except QQP, we employ a batch size of 32, while QQP uses a larger batch size of 128 due to its dataset characteristics. The model is trained uniformly for 5 epochs across all tasks with a maximum sequence length of 128. For each task, we perform a grid search over learning rates, and for each learning rate, we run three independent random experiments. For most methods, the learning rate range is $\{1\text{e-}3,8\text{e-}4,4\text{e-}4\}$ and the weight decay is set to 0.01.
@@
We fine-tune the pre-trained LLaMA3-8B model (Dubey et al. 2024) using the LLM-Adapters codebase [^3] https://github.com/AGI-Edgerunners/LLM-Adapters with evaluation via standardized lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness on the Commonsense benchmark with the Hugging Face implementation [^5] https://huggingface.co/datasets/fw407/Commonsense-15K. The fine-tuning process employs consistent hyperparameters across the AdamW optimizer. Specifically, we train for 5 epochs with a total of 588 training steps, including 20 warm-up steps. The batch size is set to 128, and the maximum sequence length is 512. We use a learning rate of $4\text{e-}4$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We evaluate the fine-tuned model on eight commonsense reasoning tasks: BoolQ (Clark et al. 2019), PIQA (Bisk et al. 2019), SIQA (Bisk et al. 2019), HellaSwag (Zellers et al. 2019), WinoGrande (Sakaguchi et al. 2019), ARC-e and ARC-c (Clark et al. 2018), and OpenbookQA (Mihaylov et al. 2018).
@@
We fine-tune the pre-trained Gemma-7B model (Gemma Team et al. 2024) using the LLM-Adapters codebase [^6] https://github.com/AGI-Edgerunners/LLM-Adapters with evaluation via standardized lm-evaluation-harness [^7] https://github.com/EleutherAI/lm-evaluation-harness on the subset of the Mathematical benchmark with the Hugging Face implementation [^8] https://huggingface.co/datasets/meta-math/MetaMathQA. We fine-tune the model using the AdamW optimizer. The training runs for 2 epochs (192 total steps, with 8 warm-up steps), using a batch size of 128 and a maximum sequence length of 512. We perform a hyperparameter search over the learning rate from the set $\{2\text{e-}4,4\text{e-}4,6\text{e-}4\}$, while keeping the optimizer parameters fixed at $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is evaluated on six mathematical reasoning benchmarks: GSM8K (Cobbe et al. 2021), MultiArith (Roy and Roth 2016), AQuA (Ling et al. 2017), SVAMP (Patel, Bhattamishra, and Goyal 2021), AddSub (Hosseini et al. 2014), and SingleEq (Koncel-Kedziorski et al. 2015).

### v2

@@ Introduction
Large-scale Pre-trained Models constitute the bedrock of modern artificial intelligence, yet their substantial fine-tuning costs drive research into Parameter-Efficient Fine-Tuning (PEFT) (ding2023parameter). PEFT techniques adapt these models with minimal computational and storage overhead through diverse strategies, ranging from inserting small, trainable adapter modules into the model’s architecture (houlsby2019parameter) to methods like prompt-tuning, which only modify the input embeddings (Lester2021ThePO; li2021prefix). Among these, LoRA (Hu2021LoRALA) is one of the most widely adopted PEFT techniques, which operates by adding a low-rank update $\mathbf{BA}$ to the pre-trained weights $\mathbf{W_{\text{pre}}}$ (i.e., $\mathbf{W=W_{pre}}+s\cdot\mathbf{BA}$).
The simplicity and effectiveness of LoRA inspire a continuous stream of research, leading to numerous variants that improve upon the original method along several key axes. These advancements primarily focus on: (i) dynamic rank allocation during training, as exemplified by AdaLoRA (Zhang2023AdaLoRAAB) and ReLORA (Lialin2023ReLoRAHT); (ii) improved initialization of low-rank matrices from pre-trained weights, proposed in methods like PiSSA (Meng2024PiSSAPS) and LoRA-GA (Wang2024LoRAGALA); and (iii) refined optimization strategies, such as those introduced in LoRA+ (Hayou2024LoRAEL) and LoRA-RITE (yen2024lora).
Our work connects to the thread of principled weight preparation, with a focus on DoRA (Weight-Decomposed Low-Rank Adaptation) (Liu2024DoRAWL). As an improvement upon LoRA, DoRA pushes this philosophy further by decoupling a pre-trained weight matrix $\mathbf{W}$ into its magnitude and direction. It then applies the low-rank update exclusively to the directional component, leaving the magnitude fixed. This process is formulated as $\mathbf{W}=\mathbf{m}\frac{\mathbf{W_{\text{pre}}}+s\mathbf{BA}}{\|\mathbf{W_{\text{pre}}}+s\mathbf{BA}\|_{c}}$.
@@
To answer this question, we conduct an in-depth empirical analysis. We find that while prior work points to stable rank (Lion2025PoLARPL), its explanatory power is limited in many practical tasks. Instead, we find that DoRA consistently increases the Singular Value Entropy of the weight update matrix. A higher entropy signifies a broader distribution of update energy across features, which prevents over-specialization and emulates the desirable behavior of full fine-tuning (Liu2025MuonIS). This suggests DoRA’s success stems from indirectly optimizing the weight update’s singular value spectrum. To validate this and address DoRA’s inefficiency, we reframe it as a more efficient Weight Conditioning method (Saratchandran2024WeightCF). This perspective reveals DoRA’s mechanism as modulating the main weight matrix with a lightweight, learnable matrix. This discovery opens a broader design space, for which we propose a unified weight conditioning framework to create more powerful and efficient PEFT methods.
Within this framework, we introduce two novel methods: (i) Pre-Diag which efficiently calibrates pre-trained weights to improve performance while reducing training time, and (ii) Skewed Orthogonal Rotation Adaptation (SORA), a powerful, norm-preserving transformation that uses parameter-efficient orthogonal rotations to align the feature space, building on Orthogonal Fine-Tuning (OFT) (Qiu2023ControllingTD).
@@
To understand the performance discrepancies among different fine-tuning methods, we first analyze the stable rank of the parameter update matrix, $\Delta\mathbf{W}=\mathbf{W}-\mathbf{W}_{\text{pre}}$, where $\mathbf{W}$ represents the final weights and $\mathbf{W_{\text{pre}}}$ the initial pre-trained weights. The stable rank is defined as $\|\Delta\mathbf{W}\|_{F}^{2}/\|\Delta\mathbf{W}\|_{2}^{2}$ (rudelson2007sampling). Our choice of this metric is motivated by prior work that has linked it to the diversity of update directions (Lion2025PoLARPL) and the expressive capacity of low-rank models (Zeng2023TheEP; Kalajdzievski2023ARS). Our initial hypothesis was that the stable rank could quantitatively explain why methods like Full Fine-Tuning (FFT) and DoRA outperform LoRA.
@@
The limitations of stable rank motivate our search for a more discerning metric. We therefore employ singular value entropy, defined as the Shannon entropy of the normalized singular value spectrum: $H(\sigma)=-\sum_{i}p_{i}\log p_{i}$, where $p_{i}=\sigma_{i}^{2}/\sum_{j}\sigma_{j}^{2}$  (7098875; doi:10.1073/pnas.97.18.10101). Intuitively, this metric measures how uniformly the update’s energy is distributed across its singular directions. A higher entropy indicates that the update is more uniformly distributed, rather than being dominated by a few high-magnitude singular values. This uniformity prevents the update from concentrating on only a few dominant features.
@@
This finding raises a crucial question: how does the DoRA formulation mechanically increase singular value entropy? The original DoRA update rule (Liu2024DoRAWL) decomposes the weight matrix $\mathbf{W}$ into a learnable magnitude $\mathbf{m}$ and a direction component:
@@
We can algebraically reformulate this operation as a form of weight conditioning (Saratchandran2024WeightCF). This reformulation reveals the effective weight change $\Delta\mathbf{W}_{\text{DoRA}}=\mathbf{W}-\mathbf{W_{\text{pre}}}$ to be:
@@
This higher-rank component is key. For singular value indices $i>r$, the LoRA update is zero (i.e., $\sigma_{i}(\Delta\mathbf{W}_{\text{LoRA}})=0$). In contrast, Weyl’s theorem (horn2012matrix) shows that DoRA maintains non-zero singular values in these dimensions:
@@
LoRA (Hu2021LoRALA) $\mathbf{W_{\text{pre}}+BA}$ Identity Matrix N/A Low-Rank Adaption
DoRA (Liu2024DoRAWL) $\mathbf{(W_{\text{pre}}+BA)D}$ Diagonal Matrix Post Weight-Decomposed
@@
This approximation makes $\mathbf{P}$ near-orthogonal while drastically reducing computational demands. (Singla2021SkewOC) first introduced the skew-orthogonal strategy into CNNs and provided a bounding for the approximation of matrix exponentials. Similarly, in SORA, for the orthogonal matrix $\mathbf{P}$, we have the following theorem:
@@
To test general language understanding, we fine-tuned DeBERTaV3-Base (He2020DeBERTaDB) on the GLUE benchmark(Wang2018GLUEAM). We test three learning rates ($\{1\text{e-}3,8\text{e-}4,4\text{e-}4\}$), each with three random seeds for robustness. Table 2 shows aggregated results comparing our Pre-Diag and SORA with LoRA and DoRA, reporting the best average performance. The findings indicate that our methods achieve superior average performance on the benchmark.
@@
For commonsense reasoning, we conduct several experiments. First, we train LLaMA3-8B (Dubey2024TheL3) on the Commonsense15k dataset and test its performance on a commonsense benchmark. The experimental results are presented in Table 3. SORA achieves the optimal average score across all tasks with different ranks, which effectively confirms the effectiveness of SORA.
@@
For mathematical reasoning, we conduct several experiments. First, we train Gemma-7B (gemma_2024) on the MetaMathQA14k dataset (yu2023metamath) and test its performance on a mathematical benchmark. The experimental results are presented in Table 4. SORA achieves the optimal average score across all tasks, which effectively confirms the effectiveness of SORA.
@@
Parameter-Efficient Fine-Tuning (PEFT) methods reduce the significant computational and storage costs associated with adapting large pre-trained models. Among various approaches, such as adapter modules (houlsby2019parameter) and prompt-tuning (Lester2021ThePO), methods based on low-rank adaptation are particularly prominent for their effectiveness and efficiency.
@@
LoRA (Hu2021LoRALA) is a cornerstone of this approach. It approximates weight updates as the product of two low-rank matrices, based on the assumption that these updates occupy a low-dimensional manifold. This method is also provably capable of approximating any target transformer model under mild assumptions (Zeng2023TheEP). This elegant mechanism inspires a wave of research that unfolds along three main threads.
@@
The first thread focuses on expressiveness and optimisation. Methods like AdaLoRA (Zhang2023AdaLoRAAB) and ReLoRA (Lialin2023ReLoRAHT) replace LoRA’s static rank with dynamic allocation schedules to place capacity where it is most needed. Others enrich the update structure beyond a simple matrix product; for instance, BTT (Potapczynski2024SearchingFE) and CoMERA (Yang2024CoMERACA) introduce tensor factorisations while maintaining linear complexity. The training process itself is also refined. LoRA+ (Hayou2024LoRAEL) advocates for distinct learning rates for the two factors, while LORO (LORO_iclr2025) and ScaleAdam (Zhang2024RiemannianPL) reframe the optimisation on Riemannian manifolds to align updates with the parameter space’s intrinsic curvature.
@@
A second thread pursues efficiency under tight resource constraints. QLoRA (Dettmers2023QLoRAEF) enables fine-tuning on a single consumer GPU by quantising the model backbone to 4-bits while keeping LoRA activations in higher precision. Complementary efforts further shrink memory footprints by compressing the adapters or sharing low-rank factors across layers and tasks, as seen in methods like LoQT and VeRA (Loeschcke2024LoQTLA; Kopiczko2023VeRAVR).
@@
The third thread involves the principled initialisation of weights. PiSSA (Meng2024PiSSAPS) and OLoRA (Bykakyz2024OLoRAOL) initialise the low-rank factors using singular vectors or orthogonal bases from the pre-trained weights, aligning the initial update with salient directions. DoRA (Liu2024DoRAWL) extends this by decomposing each weight into magnitude and direction, applying LoRA only to the directional component. Such targeted adaptation yields significant gains and highlights the value of structuring the update, not just constraining its rank.
@@
Beyond LoRA-centric innovations, other structured PEFT methods explore different constraints. One class of methods, including HRA (Yuan2024BridgingTG) and HOFT (Arcas2025HOFTHO), constructs complex orthogonal matrices from simple low-rank operations like Householder transformations. Another evolutionary line of structured methods seeks to balance expressive power, cost, and efficiency. This line evolves from the restrictive block-diagonal form of OFT (Qiu2023ControllingTD), to the denser but more expensive butterfly structure of BOFT, and finally to the efficient group-and-shuffle structure of GSOFT (Gorbunov2024GroupAS). Our method, SORA, also operates in this space. Inspired by the Cayley transform (Trockman2021OrthogonalizingCL), SORA constructs orthogonal matrices using a first-order Taylor expansion of the matrix exponential, an approximation also explored in (Singla2021SkewOC). As a future direction, SORA can be integrated with orthogonalisation techniques to enhance performance.
@@ Acknowledgments
This research was supported by Guangdong S&T Programme (No. 2024B0101010003). The computing resources were provided by Pengcheng Cloud Brain.
@@ Appendix B Appendix 2: Proof of Theorem 2
See the proof of [(Singla2021SkewOC), Theorem 3].
@@
where $(\circ)$ is due to the use of the triangle inequality; $(\star)$ is due to $\|\mathbf{A}\mathbf{B}\|_{F}\leq\|\mathbf{A}\|_{F}\|\mathbf{B}\|_{F}$ for all $\mathbf{A}\in\mathbb{R}^{m\times n},\mathbf{B}\in\mathbb{R}^{n\times m}$.
@@ Appendix C Appendix 3: Experimental Details
We fine-tune the pre-trained DeBERTa-Base model(He2020DeBERTaDB) on the GLUE benchmark using the Hugging Face implementation [^1] https://huggingface.co/transformers/model˙doc/roberta.html [^2] https://huggingface.co/datasets/nyu-mll/glue. For all tasks except QQP, we employ a batch size of 32, while QQP uses a larger batch size of 128 due to its dataset characteristics. The model is trained uniformly for 5 epochs across all tasks with a maximum sequence length of 128. For each task, we perform a grid search over learning rates, and for each learning rate, we run three independent random experiments. For most methods, the learning rate range is $\{1\text{e-}3,8\text{e-}4,4\text{e-}4\}$ and the weight decay is set to 0.01.
@@
We fine-tune the pre-trained LLaMA3-8B model (Dubey2024TheL3) using the LLM-Adapters codebase [^3] https://github.com/AGI-Edgerunners/LLM-Adapters with evaluation via standardized lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness on the Commonsense benchmark with the Hugging Face implementation [^5] https://huggingface.co/datasets/fw407/Commonsense-15K. The fine-tuning process employs consistent hyperparameters across the AdamW optimizer. Specifically, we train for 5 epochs with a total of 588 training steps, including 20 warm-up steps. The batch size is set to 128, and the maximum sequence length is 512. We use a learning rate of $4\text{e-}4$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We evaluate the fine-tuned model on eight commonsense reasoning tasks: BoolQ (Clark2019BoolQET), PIQA (Bisk2019PIQARA), SIQA (Bisk2019PIQARA), HellaSwag (zellers2019hellaswag), WinoGrande (Sakaguchi2019WinoGrande), ARC-e and ARC-c (Clark2018ThinkYH), and OpenbookQA (Mihaylov2018CanAS).
@@
We fine-tune the pre-trained Gemma-7B model (gemma_2024) using the LLM-Adapters codebase [^6] https://github.com/AGI-Edgerunners/LLM-Adapters with evaluation via standardized lm-evaluation-harness [^7] https://github.com/EleutherAI/lm-evaluation-harness on the subset of the Mathematical benchmark with the Hugging Face implementation [^8] https://huggingface.co/datasets/meta-math/MetaMathQA. We fine-tune the model using the AdamW optimizer. The training runs for 2 epochs (192 total steps, with 8 warm-up steps), using a batch size of 128 and a maximum sequence length of 512. We perform a hyperparameter search over the learning rate from the set $\{2\text{e-}4,4\text{e-}4,6\text{e-}4\}$, while keeping the optimizer parameters fixed at $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is evaluated on six mathematical reasoning benchmarks: GSM8K (cobbe2021training), MultiArith (roy2016solving), AQuA (ling2017program), SVAMP (patel2021nlp), AddSub (hosseini2014learning), and SingleEq (koncel2015parsing).